# Quantum-centric Supercomputing for Materials Science: A Perspective on Challenges and Future Directions

### v1

@@ Abstract
Computational models are an essential tool for the design, characterization, and discovery of novel materials. Hard computational tasks in materials science stretch the limits of existing high-performance supercomputing centers, consuming much of their simulation, analysis, and data resources. Quantum computing, on the other hand, is an emerging technology with the potential to accelerate many of the computational tasks needed for materials science. In order to do that, the quantum technology must interact with conventional high-performance computing in several ways: approximate results validation, identification of hard problems, and synergies in quantum-centric supercomputing. In this paper, we provide a perspective on how quantum-centric supercomputing can help address critical computational problems in materials science, the challenges to face in order to solve representative use cases, and new suggested directions.
@@
Currently, high-performance supercomputing centers dedicate a substantial amount of resources to computations in the materials science space [2, 3, 4, 5]. A fundamental issue is that ab-initio computations of materials, in the full configuration interation (CI) limit, suffer from an exponential or factorial growth of the computational resources in the system size [6, 7, 8, 9, 10, 11]. Therefore, memory and processing times requirements rapidly become untractable, as the size of the quantum systems increase. To tame this dire scaling, various approximations are employed to reduce classical computational costs, although ultimately even these either fail - for highly correlated materials - or become intractable for sufficiently large and general systems [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], thus consuming a great deal of computational resources of supercomputing centers.
Quantum computers present an attractive alternative, since many quantum algorithms avoid incurring in the exponential memory overheads of classical computations of quantum matter. For practical use-cases, quantum computers may be expected to operate embedded in classical high-performance computing (HPC) environments, so-called quantum-centric supercomputing. Intensive classical processing is required before, after, and concurrently with these quantum computations, for important reasons: alleviating the workload of quantum computers, integrate into existing classical HPC algorithms, extract signal from noisy quantum devices and exploit fault-tolerant codes to the best extent.
In this paper, we motivate that quantum-centric supercomputing is critical for materials science research and industrial development, we identify the challenges ahead to achieve practical quantum advantage on these use cases and propose some directions to address them. It is important to note that this work is intended as a perspective and not as an exhaustive review of the existing literature. More specifically, our discussion will be centered around addressing these particular aspects:
@@
Implications on the design of quantum-centric supercomputing architectures. We discuss the requirements that arise for quantum-centric supercomputing architectures. This includes evaluating the computational and operational demands, scalability, and integration hurdles of deploying these algorithms in combined quantum-classical HPC environments.
@@
    the use case represent an interesting problem in materials science.
@@
We organize the discussion as follows. Section I highlights some key existing algorithms relevant to materials science. Section II exposes classical computations in quantum-HPC workflows, keeping key algorithms and use cases as a reference. Section III goes into the details of the challenges for hybrid quantum-classical workload management and quantum-HPC integration. Section IV summarizes the state-of-the-art in classical simulation of quantum circuits, which is a necessary crucial consideration in evaluating the potential for quantum advantage. Finally, Section V gives an overview of potential use cases in materials science, bringing together considerations and threads from previous sections.
I Fundamental quantum algorithms
@@
In this section we summarize quantum algorithms that can be used for materials science applications. We consider algorithms that simulate quantum systems and mention quantum computing approaches to simulate partial differential equations (PDEs). Although the simulation of quantum systems is the most natural application of quantum computers in the materials science space, we mention quantum algorithms for PDEs since they can be useful for some well-crafted use cases. Within the simulation category, we discuss the time dynamics, ground states, open quantum systems and finite-temperature properties.
@@
Simulation of dynamics generated by a Hamiltonian of interest is one of the most natural applications for quantum computers. Quantum algorithms for approximating time evolution have been studied extensively, and we just name a few key methods here. In recent years the development of quantum signal processing culminated in an algorithm called qubitization [24], the complexity of which is measured by the number of queries to an oracle that encodes the Hamiltonian. For a target evolution time $t$ and an error tolerance $\epsilon$, qubitization makes only $O(t+\log 1/\epsilon)$ queries to the oracle, which is asymptotically optimal. However, the construction of the Hamiltonian oracle generally involves ancillary qubits, making qubitization difficult to implement in near-term and early fault-tolerant devices.
@@
In this regime, Trotterization [25, 26] (see also Sec. II.2.1), despite being the original quantum algorithm for simulating dynamics, remains one of the leading candidates for early applications, because its low overhead permits fine-grained tradeoffs between circuit depth and accuracy. Asymptotically, the gate complexity of the $p$-th order Trotterization scales as $O(t^{1+1/p}/\epsilon^{1/p})$, which can be made close to that of qubitization at large $p$. Additionally, the error of Trotterization depends on the commutativity of the terms in the Hamiltonian, allowing the algorithm to take advantage of the structure of the Hamiltonian. The final example we mention is the HHKL algorithm [27], which achieves nearly optimal scalings for Hamiltonians that are geometrically local (or even with power-law decaying interactions [28]) on some lattice. HHKL can be considered a hybrid between Trotterization and the high-accuracy methods that exploits commutation between spatially-separated Hamiltonian terms.
@@
Perhaps the most paradigmatic type of quantum simulation for materials is the approximation of ground states and low-energy states in general. While such problems are QMA-hard [29, 30] and thus believed to be intractable in worst cases, their broad utility across physics, chemistry, and materials science has made them the objects of intense study in the hope that for physically interesting cases, they may be manageable. The most heavily studied algorithm for simulating ground states in the near term is the variational quantum eigensolver (VQE, e.g., [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]), which is based on optimizing a parameterized wavefunction prepared on the quantum computer, and is well-adapted to near term due to the possibility of implementing it with low-depth circuits. Similar considerations apply to the use of finite-depth quantum approximate optimization algorithms for finding the ground state of frustrated materials [54, 55]. However, VQE has a number of drawbacks, mostly related to the difficulties of establishing convergence of the optimization and of accurately measuring the target energy to be estimated, which can result in long runtimes or getting stuck in local minima.
@@
In the near term, quantum Krylov algorithms [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] and related techniques [81] have recently emerged as a promising family of ground state simulation methods since they fall somewhere between VQE and QPE: many variants [63, 64, 67, 68, 69, 82, 71, 72, 73, 75, 77, 76, 78, 79] have provable convergence [82, 78, 77, 80] subject to similar assumptions to QPE, but the circuits they involve are only moderately more complex than those in VQE.
@@
There are several proposals for quantum algorithms for PDEs, usually based on the quantum linear systems algorithm [114]. Boundary value problems analyzed via finite element methods often result in large linear systems of equations [115]. Their matrix elements can frequently be derived analytically which may allow sidestepping the need for a QRAM. Existing literature has shown that quantum algorithms for finite element methods exhibit polynomial speedups over their classical counterparts [116]. These speedups have been studied for applications in solving the heat equation [117] and electromagnetic structure problems [118].
@@
II Classical information processing in quantum-classical workflows
II.1 Mapping to qubits
Introduction - Simulation of quantum systems is a promising use case for both noisy and fault-tolerant quantum computers. In order to simulate a physical quantum system, the governing physical Hamiltonian, for example, that of a fermionic or bosonic system, must be appropriately mapped to a qubit Hamiltonian such that the quantum computer faithfully simulates the physical system of interest. Note an appropriate mapping preserves the operator algebra: the physical operators that comprise the physical Hamiltonian and the qubit operators that comprise the qubit Hamiltonian must obey the same algebra. Fermions and bosons follow their respective algebras, commonly known as canonical anticommutation relations (CAR) and canonical commutation relations (CCR). Qubits are the so-called hard-core bosons [125].
@@ II.1 Mapping to qubits
Fermionic wavefunctions, on the other hand, are antisymmetric, which means that in second-quantization their creation and annihilation operators anticommute on different modes. Since spatially disjoint qubit operators commute, mapping fermions to qubits requires care. Also known as bosonization in the physics community (in the sense of qubits sharing the commutation relations of bosons), a collection of qubit operators can be used to embed a single fermionic creation or annihilation operator to preserve the fermionic algebra between the embedded fermionic operators. The first such mapping was the Jordan-Wigner (JW) transformation [133], and more recently this area has seen extensive research [134, 135, 136, 137, 138, 139, 140, 50, 141, 142, 143, 144, 53], with each of these references discussing new fermion-to-qubit mappings.
@@
Some of these alternative mappings may be related to the JW transformation by a unitary since conjugation by a unitary preserves the entire algebra. Others, especially those designed for particular fermionic lattices, may preserve the algebra between the relevant operators only, implied by the given lattice. A subset of unitaries, such as Cliffords or linear-reversible circuits, can be used to describe other well-known mappings including Bravyi-Kitaev (BK) [134] or parity.
@@
The art of finding and choosing a good mapping is centered around the computational resource savings. This is because different mappings can result in vastly different quantum circuits, e.g. with respect to the number of qubits or the number of quantum gates, when combined with both problem-specific and problem-agnostic circuit optimization techniques. In the near term, circuits have limited width and depth due to noise, so providing a mapping that is efficient in both of these respects is crucial to successful execution. Indeed, while the parity transform has been used to reduce the qubit counts [137] and the BK transform has been shown to admit the asymptotically optimal operator locality for generic fermionic Hamiltonians [134], customized mappings that take advantage of the input problem specifics while working in concert with circuit compilation tools such as [145, 48] have been reported to result in substantially smaller circuits [50, 53].
@@
Other mappings that aim to exploit limited connectivity in fermionic lattices to localize qubit operators over the limited qubit connectivity offered by some quantum computers have also been explored [50, 141, 142, 146, 144, 147]. With very few exceptions (e.g., [143]), most fermion-to-qubit mappings constitute mappings from products of Majorana operators to Pauli operators. Many interesting and useful configurations have been deployed within this framework. The space of mappings of this type is well defined [144], although not necessarily easy to search over. In this sense future work in constructing new useful mappings of this type constitutes a collective exercise in combinatorial optimization, which remains a useful and interesting activity, especially in the near-term regime where these details of efficiency can make or break an algorithm. The broader challenge for the field is finding ways to push beyond this paradigm, to either find new kinds of useful mappings that are not somehow equivalent to mapping Majoranas to Paulis, or to otherwise leverage our understanding of the landscape of mappings to some novel useful effect.
II.2 Hamiltonian encodings for dynamics simulation
@@
The intractability of simulating dynamics of generic quantum systems using known classical methods has driven the development of quantum computation since its inception [148], and it is the branch of quantum simulation that is most strongly believed to offer a quantum advantage [25]. Hamiltonian dynamics, or the simulation of time evolution operator $e^{-iHt}$ for a given Hamiltonian $H$, is a core component in many quantum algorithms. Generally, the Hamiltonian is represented as a sum, $H=\sum_{i}H_{i}$, of operators diagonalizable using simple unitary transformations $H_{i}=U_{i}D_{i}U_{i}^{\dagger}$ (e.g., exact squares of one-electron fermionic operators for quantum chemistry calculations, weighted sum of fully commuting Pauli operators, or one-sparse operators) whose dynamics admit more straightforward simulation. The choice of these operators and their ordering, as well as the analysis of their sparsity and commutation relations, which is usually done classically, can greatly impact the efficiency of the algorithm [149]. Depending on the decomposition, a variety of well-known techniques, such as Trotterization and linear combination of unitaries, can then be used to compile the dynamics operator into fundamental gate operations.
II.2.1 Trotterization
@@
The basic idea of Trotterization is to break down the evolution of a given quantum system into a series of smaller, more manageable steps. For a given Hamiltonian operator $H=\sum_{i}H_{i}$, the first-order Trotter approximation is given by the product formula
@@
Trotterization introduces errors in the simulation as shown by the above examples. At the highest level, this error depends on the order $p$ of the product formula used and scales asymptotically as $O(t^{p+1}/n^{p})$ for $n$ Trotter steps [25, 153]. However, this error is coming from non-commuting terms in the Hamiltonian decomposition, which suggests that tighter error bounds may be available that take advantage of the specific structure of commutation. In addition, in quantum computation, Trotterized time evolutions are always applied to some state, in which case we would ultimately care about the error in the resulting approximate time-evolved state rather than the operator error in the Trotter approximation itself. Research into providing tighter bounds and optimizing this error (e.g. by grouping and reordering of terms) is important and ongoing [154, 26, 149].
@@
for Pauli operators $P_{i}$ and real coefficients $c_{i}$, an upper bound on the number of steps can be given in terms of the $l1$-norm of the coefficients $|||H|||=\sum_{i}|c_{i}|$. Such bounds, however, ignore commutativity between terms and tend to grossly overestimate the number of Trotter steps required to approximate the Hamiltonian evolution to a certain error. Indeed, in the trivial case where all terms in the Hamiltonian representation commute pairwise, a single Trotter step is enough to represent the Hamiltonian evolution exactly. A recent study on the theory of Trotter errors [26] provided more accurate bounds on the number of steps for a $p$-th order formula in terms of norms of nested commutators $\sum_{\imath_{1}\cdots i_{p+1}}||[H_{i_{p+1}},[\cdots[H_{i_{2}},H_{i_{1}}]]\cdots]||$. Computing the precise commutator-scaling is exceptionally useful for reducing the number of Trotter steps, and, consequently, the quantum resources required, without sacrificing accuracy. It is, however, computationally expensive, as the computational cost, $O(N^{p+1})$, scales polynomially with the number of terms $N$ in the Hamiltonian. Classical workflows (e.g. ordering of the Hamiltonian terms to reduce circuit depth) and their efficient implementation (e.g. leveraging high-performance parallel computing for accurate commutator-scaling evaluation) will play an important role in reducing the quantum resources required for practical applications and shortening the timeline to practical quantum advantage.
@@
One final point we would like to make here is that the discussion above focuses on generic Hamiltonian simulation and the analysis of the Hamiltonian operator. In practice, especially when targeting early quantum applications, one is interested in the accurate evolution of a given Hamiltonian, for a given time, with a given initial quantum state. The details of choosing the appropriate Trotterization (and, more generally speaking, the Hamiltonian simulation approach) will be very much application-specific, and, when compared to results obtained by only analyzing the Hamiltonian operator, further reduction of the estimated resources will most certainly be achieved.
II.2.2 Linear Combination of Unitaries

### v2

@@ Abstract
Computational models are an essential tool for the design, characterization, and discovery of novel materials. Computationally hard tasks in materials science stretch the limits of existing high-performance supercomputing centers, consuming much of their resources for simulation, analysis, and data processing. Quantum computing, on the other hand, is an emerging technology with the potential to accelerate many of the computational tasks needed for materials science. In order to do that, the quantum technology must interact with conventional high-performance computing in several ways: approximate results validation, identification of hard problems, and synergies in quantum-centric supercomputing. In this paper, we provide a perspective on how quantum-centric supercomputing can help address critical computational problems in materials science, the challenges to face in order to solve representative use cases, and new suggested directions.
I Introduction
@@
Currently, high-performance supercomputing centers dedicate a substantial amount of resources to computations in the materials science space [2, 3, 4, 5]. A fundamental issue is that ab-initio computations of materials, in the full configuration interation (CI) limit, suffer from an exponential or factorial growth of the computational resources in the system size [6, 7, 8, 9, 10, 11]. Therefore, memory and processing times requirements rapidly become intractable, as the size of the quantum systems increase. To tame this dire scaling, various approximations are employed to reduce classical computational costs, however, even these either fail – for highly correlated materials – or become intractable for sufficiently large and general systems [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Quantum computers present an attractive alternative, since many quantum algorithms avoid incurring in the exponential memory overheads of classical computations of quantum matter. For practical use cases, quantum computers may be expected to operate embedded in classical high performance computing (HPC) environments, so-called Quantum-centric Supercomputing (QCSC). Intensive classical processing is required before, after, and concurrently with these quantum computations, for a number of reasons: alleviating the workload of quantum computers, integration into existing classical HPC algorithms, signal extraction from noisy quantum devices, and exploiting fault-tolerant codes to the best extent.
In this paper, we motivate the need for quantum-centric supercomputing in materials science research and industrial development. We also identify the challenges ahead to achieve practical quantum advantage for materials science use cases and propose directions to address them. It is important to note that this work is intended as a perspective and not as an exhaustive review of the existing literature. More specifically, our discussion will be centered around addressing several particular aspects:
@@
Implications on the design of quantum-centric supercomputing architectures. We discuss the requirements that arise for quantum-centric supercomputing architectures. This includes evaluating the computational and operational demands, scalability, and integration hurdles when deploying these algorithms in coupled quantum-classical HPC environments.
@@
    the use case represents an interesting problem in materials science.
@@
Our discussion is organized as follows. Section II highlights some key existing algorithms relevant to materials science. Section III exposes classical computations in quantum-HPC workflows, keeping key algorithms and use cases as a reference. In Sec.tion IV, we detail of the challenges for hybrid quantum-classical workload management and quantum-HPC integration, and in Sec. V give a perspective on the main components of a programming model for such systems. Section VI summarizes the state-of-the-art in classical simulation of quantum circuits, which is a necessary crucial consideration in evaluating the potential for quantum advantage. Finally, Sec. VII gives an overview of potential use cases in materials science, bringing together considerations and threads from previous sections.
II Fundamental quantum algorithms
@@
In this section we summarize quantum algorithms that can be used for materials science applications. We consider algorithms that simulate quantum systems and mention quantum computing approaches to simulate partial differential equation (PDE). Although the simulation of quantum systems is the most natural application of quantum computers in the materials science space, we mention quantum algorithms for PDEs since they can be useful for some well-crafted use cases. Within the simulation category, we discuss the time dynamics, ground states, open quantum systems and finite-temperature properties.
@@
Simulation of dynamics generated by a Hamiltonian of interest is one of the most natural applications for quantum computers. Quantum algorithms for approximating time evolution have been studied extensively, and so here briefly describe a few key methods. In recent years the development of quantum signal processing culminated in an algorithm called qubitization [24], the complexity of which is measured by the number of queries to an oracle that encodes the Hamiltonian. For a target evolution time $t$ and an error tolerance $\epsilon$, qubitization makes only $O(t+\log 1/\epsilon)$ queries to the oracle, which is asymptotically optimal. However, the construction of the Hamiltonian oracle generally involves ancillary qubits, making qubitization difficult to implement in near-term and early fault-tolerant devices.
@@
In this regime, Trotterization [25, 26] (see also Sec. III.2.1), the original quantum algorithm for simulating dynamics, remains one of the leading candidates for early applications, because its low overhead permits fine-grained tradeoffs between circuit depth and accuracy. Asymptotically, the gate complexity of the $p$-th order Trotterization scales as $O(t^{1+1/p}/\epsilon^{1/p})$, which can be made close to that of qubitization at large $p$. Additionally, the error of Trotterization depends on the commutativity of the terms in the Hamiltonian, allowing the algorithm to take advantage of the structure of the Hamiltonian. The final example we mention is the HHKL algorithm [27], which achieves nearly optimal scalings for Hamiltonians that are geometrically local (or even with power-law decaying interactions [28]) on some lattice. HHKL can be considered a hybrid between Trotterization and the high-accuracy methods that exploits commutation between spatially-separated Hamiltonian terms.
@@
Perhaps the most paradigmatic type of quantum simulation for materials is the approximation of ground states and low-energy states in general. While such problems are Quantum Merlin Arthur (QMA)-hard [29, 30] and thus believed to be intractable in worst cases, their broad utility across physics, chemistry, and materials science has made them the objects of intense study in the hope that for physically interesting cases, they may be manageable. The most heavily studied algorithm for simulating ground states in the near term is the Variational Quantum Eigensolver (VQE), e.g., [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]), which is based on optimizing a parameterized wavefunction prepared on the quantum computer, and is well-adapted to near term due to the possibility of implementing it with low-depth circuits. Similar considerations apply to the use of finite-depth quantum approximate optimization algorithms for finding the ground state of frustrated materials [54, 55]. However, VQE has a number of drawbacks, mostly related to the difficulties of establishing convergence of the optimization and of accurately measuring the target energy to be estimated, which can result in long runtimes or getting stuck in local minima.
@@
In the near term, quantum Krylov algorithms [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81] and related techniques [82] have recently emerged as a promising family of ground state simulation methods since they fall somewhere between VQE and QPE: many variants [63, 64, 67, 68, 69, 83, 71, 72, 73, 75, 77, 76, 78, 79, 80] have provable convergence [83, 80, 78, 77, 81] subject to similar assumptions to QPE, but the circuits they involve are only moderately more complex than those in VQE.
@@
There are several proposals for quantum algorithms for PDEs, usually based on the quantum linear systems algorithm [115]. Boundary value problems analyzed via finite element methods often result in large linear systems of equations [116]. Their matrix elements can frequently be derived analytically which may allow sidestepping the need for a quantum random access memory (QRAM). Existing literature has shown that quantum algorithms for finite element methods exhibit polynomial speedups over their classical counterparts [117]. These speedups have been studied for applications in solving the heat equation [118] and electromagnetic structure problems [119].
@@
III Classical information processing in quantum-classical workflows
III.1 Mapping to qubits
Simulation of quantum systems is a promising use case for both noisy and fault-tolerant quantum computers. In order to simulate a physical quantum system, the governing physical Hamiltonian, for example, that of a fermionic or bosonic system, must be appropriately mapped to a qubit Hamiltonian such that the quantum computer faithfully simulates the physical system of interest. Note an appropriate mapping preserves the operator algebra: the physical operators that comprise the physical Hamiltonian and the qubit operators that comprise the qubit Hamiltonian must obey the same algebra. Fermions and bosons follow their respective algebras, commonly known as canonical canonical anticommutation relations (CAR) and canonical commutation relations (CCR). Qubits are the so-called hard-core bosons [126].
@@ III.1 Mapping to qubits
Fermionic wavefunctions, on the other hand, are antisymmetric, which means that in second-quantization their creation and annihilation operators anticommute on different modes. Since spatially disjoint qubit operators commute, mapping fermions to qubits requires care. Also known as bosonization in the physics community (in the sense of qubits sharing the commutation relations of bosons), a collection of qubit operators can be used to embed a single fermionic creation or annihilation operator to preserve the fermionic algebra between the embedded fermionic operators. The first such mapping is the Jordan-Wigner (JW) transformation [134], and more recently this area has seen extensive research [135, 136, 137, 138, 139, 140, 141, 50, 142, 143, 144, 145, 53], with each of these references discussing new fermion-to-qubit mappings.
@@
Some of alternative mappings may be related to the JW transformation by a unitary transformation since conjugation by a unitary preserves the entire algebra. Others, especially those designed for particular fermionic lattices, may preserve the algebra between the relevant operators only, implied by the given lattice. A subset of unitaries, such as Cliffords or linear-reversible circuits, can be used to describe other well-known mappings including Bravyi-Kitaev (BK) [135] or parity.
@@
The art of finding and choosing a good mapping is centered around the computational resource savings. This is because different mappings can result in vastly different quantum circuits; e.g., with respect to the number of qubits or the number of quantum gates, when combined with both problem-specific and problem-agnostic circuit optimization techniques. In the near term, circuits will necessarily have limited width and depth due to noise, so providing a mapping that is efficient in both dimensions is crucial for reliable execution. Indeed, while the parity transform has been used to reduce the qubit counts [138] and the BK transform has been shown to admit the asymptotically optimal operator locality for generic fermionic Hamiltonians [135], customized mappings that take advantage of the input problem specifics while working in concert with circuit compilation tools, e.g.,  [146, 48], have been reported to result in substantially smaller circuits [50, 53].
@@
Other mappings that aim to exploit limited connectivity in fermionic lattices to localize qubit operators over the limited qubit connectivity offered by some quantum computers have also been explored [50, 142, 143, 147, 145, 148]. With very few exceptions, e.g., [144], most fermion-to-qubit mappings constitute mappings from products of Majorana operators to Pauli operators. Many interesting and useful configurations have been deployed within this framework. The space of mappings of this type is well defined [145], although not necessarily easy to search over. In this sense future work in constructing new useful mappings of this type constitutes a collective exercise in combinatorial optimization, which remains a useful and interesting activity, especially in the near-term regime where these details of efficiency can make or break an algorithm. The broader challenge for the field is finding ways to push beyond this paradigm, to either find new kinds of useful mappings that are not somehow equivalent to mapping Majoranas to Paulis, or to otherwise leverage our understanding of the landscape of mappings to some novel useful effect.
III.2 Hamiltonian encodings for dynamics simulation
@@
The intractability of simulating dynamics of generic quantum systems using known classical methods has driven the development of quantum computation since its inception [149], and it is the branch of quantum simulation that is most strongly believed to offer a quantum advantage [25]. Hamiltonian dynamics, or the simulation of time evolution operator, $e^{-iHt}$, for a given Hamiltonian, $H$, is a core component in many quantum algorithms. Generally, the Hamiltonian is represented as a sum, $H=\sum_{i}H_{i}$, of operators which are diagonalizable using simple unitary transformations, $H_{i}=U_{i}D_{i}U_{i}^{\dagger}$; e.g., exact squares of one-electron fermionic operators for quantum chemistry calculations, weighted sum of fully commuting Pauli operators, or one-sparse operators, whose dynamics admit more straightforward simulation. The choice of these operators and their ordering, as well as the analysis of their sparsity and commutation relations, which is usually done classically, can greatly impact the efficiency of the algorithm [150]. Depending on the decomposition, a variety of well-known techniques, such as Trotterization and linear combination of unitaries, can then be used to compile the dynamics operator into fundamental gate operations.
III.2.1 Trotterization
@@
The basic idea of Trotterization is to break down the evolution of a given quantum system into a series of smaller, more computationally manageable steps. For a given Hamiltonian operator $H=\sum_{i}H_{i}$, the first-order Trotter approximation is given by the product formula
@@
Trotterization introduces errors in the simulation as shown by the above examples. At the highest level, this error depends on the order $p$ of the product formula used and scales asymptotically as $O(t^{p+1}/n^{p})$ for $n$ Trotter steps [25, 154]. However, this error is coming from non-commuting terms in the Hamiltonian decomposition, which suggests that tighter error bounds may be available when the specific structure of commutation is leveraged. In addition, in quantum computation, Trotterized time evolutions are always applied to some state, in which case one would ultimately care about the error in the resulting approximate time-evolved state rather than the operator error in the Trotter approximation itself. Research into providing tighter bounds and optimizing this error (e.g. by grouping and reordering of terms) is important and ongoing [155, 26, 150].
@@
for Pauli operators $P_{i}$ and real coefficients $c_{i}$, an upper bound on the number of steps can be given in terms of the $\ell_{1}$-norm of the coefficients, $|||H|||=\sum_{i}|c_{i}|$. Such bounds, however, ignore commutativity between terms and tend to grossly overestimate the number of Trotter steps required to approximate the Hamiltonian evolution to a certain error. Indeed, in the trivial case where all terms in the Hamiltonian representation commute pairwise, a single Trotter step is enough to represent the Hamiltonian evolution exactly. A recent study on the theory of Trotter errors [26] provides more accurate bounds on the number of steps for a $p$th order formula in terms of norms of nested commutators $\sum_{\imath_{1}\cdots i_{p+1}}||[H_{i_{p+1}},[\cdots[H_{i_{2}},H_{i_{1}}]]\cdots]||$. Computing the precise commutator-scaling is exceptionally useful for reducing the number of Trotter steps, and, consequently, the quantum resources required, without sacrificing accuracy. It is, however, computationally expensive, as the computational cost, $O(N^{p+1})$, scales polynomially with the number of terms $N$ in the Hamiltonian. Classical workflows (e.g. ordering of the Hamiltonian terms to reduce circuit depth) and their efficient implementation (e.g. leveraging high-performance parallel computing for accurate commutator-scaling evaluation) will play an important role in reducing the quantum resources required for practical applications and shortening the timeline to practical quantum advantage.
@@
One final point to note is that the above discussion focuses on generic Hamiltonian simulation and the analysis of the Hamiltonian operator. In practice, especially when targeting early quantum applications, one is interested in the accurate evolution of a given Hamiltonian, for a given time, with a given initial quantum state. The details of choosing the appropriate Trotterization (and, more generally speaking, the Hamiltonian simulation approach) will be very much application-specific, and, when compared to results obtained by only analyzing the Hamiltonian operator, further reduction of the estimated resources will most certainly be achieved.
III.2.2 Linear Combination of Unitaries